Show that the vectors a = 3i − j + 4k, b = i − 3j − 2k and c = 4i − 3j + 2k are
linearly independent. Find numbers α, β and γ such that d = 2i + 3j − k can be
expressed in the form d = αa + βb + γc.
1
Expert's answer
2019-10-16T10:15:05-0400

Solution: Firstly, show that these vectors are linearly independent. Write our vectors vertical in matrix. Use Gaussian elimination for linearly independent vectors:"\begin{pmatrix}
 3 & 1 & 4 \\
 -1 & -1 & -3\\
 4 & -2& 2
\end{pmatrix}"

Divide the first row by 3:

"\begin{pmatrix}
 1 & 1/3 & 4/3 \\
 -1 & -1 & -3\\
 4 & -2& 2
\end{pmatrix}"

Substitute the second and third rows from the first multiplied by -1 and 4:

"\begin{pmatrix}
 1 & 1/3 & 4/3 \\
 0 & -8/3 & -5/3\\
 0 & -10/3 & -10/3
\end{pmatrix}"

Divide the second row by 8/3:

"\begin{pmatrix}
 1 & 1/3 & 4/3 \\
 0 & -1 & -5/8\\
 0 & -10/3 & -10/3
\end{pmatrix}"

Substitute the third row from second multiplied by 10/3:

"\begin{pmatrix}
 1 & 1/3 & 4/3 \\
 0 & -1 & -5/8\\
 0 & 0 & -5/4
\end{pmatrix}"

So we have 3 rows without zero (0) in all three places. So we have linearly independent vectors.

Next task is to find unknown numbers α, β and γ for another vector. Write a supplemented matrix (from previous one) with numbers of another vector.

"\begin{pmatrix}
 3 & 1 & 4 && 2 \\
 -1 & -1 & -3 && 3\\
 4 & -2& 2 && -1
\end{pmatrix}"

Divide the first row by 3:

"\begin{pmatrix}
 1 & 1/3 & 4/3 && 2/3 \\
 -1 & -1 & -3 && 3\\
 4 & -2& 2 && -1
\end{pmatrix}"

Substitute the second and the third rows from the first multiplied by 1 and -4

"\begin{pmatrix}
 1 & 1/3 & 4/3 && 2/3 \\
 0 & -8/3 & -5/3 && 11/3\\
 0 & -10/3& -10/3 && -11/3
\end{pmatrix}"

Divide the second row by 8/3

"\begin{pmatrix}
 1 & 1/3 & 4/3 && 2/3 \\
 0 & -1 & -5/8 && 11/8\\
 0 & -10/3& -10/3 && -11/3
\end{pmatrix}"

Substitute the third row from the second multiplied by 10/3:

"\begin{pmatrix}
 1 & 1/3 & 4/3 && 2/3 \\
 0 & -1 & -5/8 && 11/8\\
 0 & 0& -5/4 && -33/4
\end{pmatrix}"

Use Gaussian elimination in another side - from top to bottom and find numbers in the last column when we would have the identity matrix.

Divide the third row by -5/4

"\begin{pmatrix}
 1 & 1/3 & 4/3 && 2/3 \\
 0 & -1 & -5/8 && 11/8\\
 0 & 0& 1 && 33/5
\end{pmatrix}"

Divide the second row by -1:

"\begin{pmatrix}
 1 & 1/3 & 4/3 && 2/3 \\
 0 & 1 & 5/8 && -11/8\\
 0 & 0& 1 && 33/5
\end{pmatrix}"

Add the third row to second multiplied by -5/8

"\begin{pmatrix}
 1 & 1/3 & 4/3 && 2/3 \\
 0 & 1 & 0 && -11/2\\
 0 & 0& 1 && 33/5
\end{pmatrix}"

Add the second row to first multiplied by -1/8

"\begin{pmatrix}
 1 & 0 & 4/3 && 5/2 \\
 0 & 1 & 0 && -11/2\\
 0 & 0& 1 && 33/5
\end{pmatrix}"

Add the third row to first multiplied by -4/3

"\begin{pmatrix}
 1 & 0 & 0 && -63/10 \\
 0 & 1 & 0 && -11/2\\
 0 & 0& 1 && 33/5
\end{pmatrix}"

We have the identity matrix and have numbers of our vector. Rewrite the system:

α = - 63/10; β = -11/2; γ = 33/5;

or

α = - 6,3; β = -5,5; γ = 6,6;

Answer: Yes, these vectors are linearly independent. Numbers of the last vector are

α = - 6,3; β = -5,5; γ = 6,6